OpenStudy (anonymous):

Find the slope of any line perpendicular to the line through (6,1) and (−4,−6).

OpenStudy (btaylor):

The slope of the line through these points is: (y1-y2)/(x1-x2) = (1-(-6))/(6-(-4)) = (1+6)/(1+4) = 7/5 The slope of a perpendicular line is the negative reciprocal of that slope, so it will be -5/7.

OpenStudy (anonymous):

its still wrong

OpenStudy (btaylor):

oops. Should be (1-(-6))/(6-(-4)) = (1+6)/(6+4) = 7/10, so the perp. slope will be -10/7. my bad. sry.

OpenStudy (anonymous):

thats ok thanks

OpenStudy (btaylor):

yw.